The radii of curvature of the faces of a double convex lens are
20 cm and 25 cm. Its focal length is 20 cm. Calculate the
refractive index of the material of the lens.

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
1/f = (n –1) [1/R1 – 1/R2]
1/20 = (n –1) [1/20 – 1/-25]
1/20 = (n –1) [9/100]
9n – 9 = 5
9n = 14
n = 14/9
n = 1.555
 
Regards
Arun (askIITians forum expert)